# Robust Estimation of Polyserial Correlation

### v1

@@ Abstract
The association between a continuous and an ordinal variable is commonly modeled through the polyserial correlation model. However, this model, which is based on a partially-latent normality assumption, may be misspecified in practice, due to, for example (but not limited to), outliers or careless responses. We demonstrate that the typically used maximum likelihood (ML) estimator is highly susceptible to such misspecification: One single observation not generated by partially-latent normality can suffice to produce arbitrarily poor estimates. As a remedy, we propose a novel estimator of the polyserial correlation model designed to be robust against the adverse effects of observations discrepant to that model. The estimator achieves robustness by implicitly downweighting such observations; the ensuing weights constitute a useful tool for pinpointing potential sources of model misspecification. We show that the proposed estimator generalizes ML and is consistent as well as asymptotically Gaussian. As price for robustness, some efficiency must be sacrificed, but substantial robustness can be gained while maintaining more than 98% of ML efficiency. We demonstrate our estimator’s robustness and practical usefulness in simulation experiments and an empirical application in personality psychology where our estimator helps identify outliers. Finally, the proposed methodology is implemented in free open-source software.
@@ 1 Introduction
The polyserial correlation model postulates the existence of a latent continuous variable that underlies and governs the observed ordinal variable through an unobserved discretization process. The correlation between the observed continuous variable and the latent variable is called polyserial correlation, whereas the correlation between the observed continuous and the observed ordinal variable is called point polyserial correlation, where the latter can be computed from the former. For identification, the polyserial correlation model assumes that the observed continuous variable and the latent continuous variable are jointly normally distributed, thereby making it a partially-latent normality model. Estimation is typically conducted by means of maximum likelihood (Cox,, 1974; Olsson et al.,, 1982). However, the validity of the partially-latent normality assumption is often questionable in practice (e.g., Barbiero,, 2025; Demirtas & Hedeker,, 2016; Bedrick,, 1995).
@@
The contributions of this paper are twofold. First, we study estimation of the polyserial correlation model when a (possibly empty) subset of the observed data are of low-quality and therefore may have not been generated by a normal distribution, such as (but not limited to) outliers in the continuous variable and/or careless responses in the ordinal variable. Consequently, such observations are uninformative for estimating the polyserial correlation model. This situation is called partial model misspecification because the assumption of partially-normality might be violated for parts of the data, but is satisfied for the remaining data. We demonstrate that already one single uninformative observation can suffice for maximum likelihood (ML) estimation of polyserial correlation to yield arbitrary results and, furthermore, that the accuracy of ML estimation is highly susceptible to even minor misspecification of partially-latent normality. Partial misspecification stems from classic literature on robust statistics (e.g., Huber & Ronchetti,, 2009) where it is known as Huber contamination model, owing to Huber, (1964).
@@
Second, in wake of the non-robustness of ML estimation to uninformative observations generated by partially-latent nonnormality, we propose an alternative estimator that is designed to be robust against such partial misspecification. The proposed estimator leverages the theory of minimum density power divergence estimation (Basu et al.,, 1998) and achieves robustness by implicitly downweighting observations that cannot be sufficiently well fitted by the polyserial model. In line with the partial misspecification framework, the estimator allows the model to be misspecified for an unknown fraction of uninformative observations, but makes no assumption on how and where partial misspecification occurs (which may be absent altogether). Consequently, partial misspecification can manifest through an unlimited and unrestricted variety of ways, such as (but not limited to) outliers or careless responses. Conversely, if the polyserial correlation model is correctly specified for all observations in a sample, then the estimate of our robust estimator asymptotically coincides with that of ML, thereby generalizing ML estimation. Moreover, we show that the robust estimator is consistent and asymptotically normally distributed, thereby allowing for statistical inference.
@@
In robust statistics, there is a well-established fundamental tension between robustness and efficiency for estimation procedures for data involving continuous variables (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986). Our estimator is no exception: The price to pay for its enhanced robustness manifests in the form of diminished efficiency compared to ML. However, we show that sacrificing as little as 2% of efficiency suffices to obtain a substantial gain in robustness, so the efficiency loss is only comparatively minor. An additional price of our robust estimator is increased computational intensity. Nevertheless, using our implementation, the robust estimator usually executes in less than five seconds on a regular laptop, so the overall computational burden should remain small in practical applications. This implementation of our proposed methodology is publicly and freely available as part of the R package robcat (for “ROBust CATegorical data analysis”; Welz et al., 2025a, ) on CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat.
@@ 2 Literature
In a broader context, violations of normality assumptions have been studied extensively in the psychometric literature, especially with respect to the distribution of latent variables (e.g., Lyhagen & Ornstein,, 2023; Monroe,, 2018; Asparouhov & Muthén,, 2016; Roscino & Pollice,, 2006; Yuan et al.,, 2004, and references therein). A particular focus of recent literature has been the polychoric correlation model of Pearson & Pearson, (1922), which models through a latent bivariate normality distribution the association of two latent variables that govern two observed ordinal variables (see Olsson,, 1979, for a modern exposition). Foldnes & Grønneberg, (2020) and Jin & Yang-Wallentin, (2017) show that ML estimation of polychoric correlation is highly susceptible to distributional misspecification, leading to possibly large biases in SEM analyses based on polychoric correlation (Foldnes & Grønneberg,, 2022; Grønneberg & Foldnes,, 2022). [^1] It should be noted that ML estimation of polychoric correlation was originally believed to be fairly robust against distributional misspecification of latent normality (e.g., Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004), an assessment based on simulation experiments using the method of Vale & Maurelli, (1983) for generating nonnormal bivariate data that are then discretized to be ordinal. However, Grønneberg & Foldnes, (2019) show that ordinal data generated in this way are indistinguishable from ordinal data generated by latent normality. Using a method that ensures proper violation of latent normality (the VITA method of Grønneberg & Foldnes,, 2017) reveals strong susceptibility of ML estimation of polychoric correlation to such violations (e.g., Foldnes & Grønneberg,, 2020, 2022). Welz et al., 2025b are concerned with partial misspecification of the polychoric correlation model, where latent normality is violated due to a (possibly empty) subset of data points that were generated by an unspecified and unknown nonnormal process. We use the same (partial) misspecification framework in this paper. Welz et al., 2025b show that having in a sample already about 5% of such observations, commonly referred to as contamination, can suffice for a substantial estimation bias. Contamination might arise due to, for instance but not limited to, careless responding to polytomous items, which has been identified as a major threat to the validity of psychometric analyses (e.g., Alfons & Welz,, 2024; Ward & Meade,, 2023; Arias et al.,, 2020; Bowling et al.,, 2016; Huang et al.,, 2015; Meade & Craig,, 2012, and references therein). As a remedy, Welz et al., 2025b propose a fully efficient contamination-robust estimator of the polychoric correlation model that makes no assumptions on the prevalence and type of contamination (which is possibly absent altogether). Their estimator exploits the theory of $C$-estimation (Welz,, 2024), being a general framework of robust estimation with categorical data.
@@ 3.1 The polyserial correlation model
where we adopt the conventions $\tau_{0}=$-$\infty$ and $\tau_{r}=+\infty$. The joint distribution function of the observed $(X,Y)$ under the polyserial model can now be expressed as
$\begin{split}{P}_{XY}\left(x,y;\bm{\theta}\right)=\mathbb{P}_{\bm{\theta}}\left[X\leq x,Y\leq y\right]&=\int_{-\infty}^{x}\sum_{\{w\in\mathcal{Y}:w\leq y\}}{p}_{XY}\left(u,w;\bm{\theta}\right)\textnormal{d}u\\ &=\int_{-\infty}^{x}\int_{-\infty}^{\tau_{y}}{p}_{X\eta}\left(u,v;\bm{\theta}\right)\textnormal{d}v\textnormal{d}u,\end{split}$ (4)
@@ 3.2 Point polyserial correlation
In addition to the correlation between the observed $X$ and the latent $\eta$, one might also be interested in the correlation between $X$ and the observed ordinal $Y$. The correlation between the observed $X$ and $Y$ is known as point polyserial correlation. In order to identify the desired point polyserial correlation $\mathrm{Cor}\left[X,\ Y\right]$, one needs to assign a numerical interpretation to the $r$ answer categories of $Y$, that is, introduce a scoring system. Recall that we assume that $Y$ takes values in the set $\mathcal{Y}=\{1,2,\dots,r\}$, where the elements in $\mathcal{Y}$ need not admit a numerical interpretation. Assigning a scoring system introduces a numeric meaning to each response category. We refer to Fernández et al., (2020), Ivanova & Berger, (2001), and Olsson et al., (1982, pp. 340–341) for discussions of different scoring systems.
$\begin{split}\mu_{Y}&=\mathbb{E}\left[Y\right]=\sum_{j=1}^{r}y_{j}\cdot{p}_{Y}\left(y_{j};\bm{\theta}\right),\qquad\textnormal{and}\\ \sigma^{2}_{Y}&=\mathbb{V}\mathrm{ar}\left[Y\right]=\sum_{j=1}^{r}y_{j}^{2}\cdot{p}_{Y}\left(y_{j};\bm{\theta}\right)-\mu_{Y}^{2}.\end{split}$ (5)
$\mu_{XY}=\mathbb{E}\left[XY\right]=\mu_{X}\left(y_{r}-\sum_{j=1}^{r-1}\Phi\left(\tau_{j}\right)\big(y_{j+1}-y_{j}\big)\right)+\rho\sigma_{X}\sum_{j=1}^{r-1}\phi\left(\tau_{j}\right)\big(y_{j+1}-y_{j}\big),$ (6)
$\widetilde{\rho}=\mathrm{Cor}\left[X,\ Y\right]=\frac{\mathrm{Cov}\left[X,\ Y\right]}{\sigma_{X}\sigma_{Y}}=\frac{1}{\sigma_{X}\sigma_{Y}}\left(\mu_{XY}-\mu_{X}\mu_{Y}\right),$ (7)
where explicit expressions for $(\mu_{Y},\sigma^{2}_{Y})$ and $\mu_{XY}$ are provided in Equations (5) and (6), respectively. The correlation coefficient $\widetilde{\rho}$ between the continuous $X$ and the ordinal $Y$ is referred to as point polyserial coefficient. For a given scoring system, $\widetilde{\rho}$ can be computed from the parameter vector $\bm{\theta}=\left(\rho,\mu,\sigma^{2},\bm{\tau}^{\top}\right)^{\top}$ of the polyserial model, a property established by Olsson et al., (1982). In particular, an estimator of $\widetilde{\rho}$ can be constructed by substituting $\bm{\theta}$ for an estimate thereof in Equations (5)–(7).
For the remainder of this paper, we resort back to the original assumption that the support of the ordinal $Y$ is given by a set of adjacent integers without numeric interpretation, $\mathcal{Y}=\{1,2,\dots,r\}$, which, we reiterate, is without loss of generality for identifying the polyserial correlation coefficient.
@@ 3.3 Maximum likelihood estimation
Alternative estimators have been proposed by Bedrick & Breslin, (1996), Lord, (1963), and Brogden, (1949), with Bedrick, (1992, 1990), Koopman, (1983) and Kraemer, (1981) studying the theoretical properties of the latter two. We discuss these approaches in more detail in Section 4.2, after conceptualizing misspecification of the polyserial model.
@@ 4.1 Partial misspecification of the polyserial model
The polyserial model is misspecified if at least one observation for the continuous-ordinal variable pair $(X,Y)$ has not been generated by the bivariate partially-latent normality model in (2). Akin to Welz et al., 2025b , we consider a partial misspecification framework where only a fraction $(1-\varepsilon)$ of observations in a given sample are generated by an underlying normal distribution $P_{X\eta}$ with true parameter $\bm{\theta}_{*}$, whereas a fixed but unknown fraction $\varepsilon$ of observations are generated from some different but unspecified underlying distribution $H_{X\eta}$. Since $H_{X\eta}$ is unspecified, its correlation structure may differ from $P_{X\eta}$ so that observations $(X,Y)$ generated by the underlying $H_{X\eta}$ (after discretization) may be uninformative for the true polyserial correlation coefficient.
@@
Misspecification models of the type in (11) are standard in the robust statistics literature, where they are known as Huber contamination models, owing to pioneering work of Huber, (1964). Following Welz et al., 2025b , we therefore adopt terminology from robust statistics and call $\varepsilon$ the contamination fraction, the uninformative $H_{X\eta}$ the contamination distribution (or simply contamination), and $F_{\varepsilon,X\eta}$ the contaminated distribution. In the case of a zero-valued contamination fraction ($\varepsilon=0$), there is no misspecification so that the polyserial model is correctly specified. Indeed, if $\varepsilon=0$, then $F_{0,X\eta}(\cdot,\cdot)={P}_{X\eta}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$ and $F_{0,XY}(\cdot,\cdot)={P}_{XY}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$, so the partial misspecification framework in (11) nests the correctly specified polyserial model.
Neither the contamination fraction $\varepsilon$ nor the contamination distribution $H_{X\eta}$ in (11) are assumed to be known. Thus, both quantities are left completely unspecified in practice, which “means that we are not making any assumptions on the degree, magnitude, or type of contamination (which is possibly absent altogether)” (Welz et al., 2025b, ). Hence, the polyserial model may be misspecified due to an unlimited variety of reasons, for instance but not limited to outliers or nonnormality in the continuous $X$ , and/or careless responding or item misunderstanding in the ordinal $Y$. Because $\varepsilon$ and $H_{X\eta}$ are unspecified in practice, the polyserial model distribution ${P}_{XY}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$ remains the distribution of interest: Our only aim is to estimate the parameter $\bm{\theta}_{*}$ of the polyserial model while reducing the impact of potential contamination in the observed data. Consequently, the contaminated distributions $F_{\varepsilon,XY}$ and $F_{\varepsilon,X\eta}$ are never estimated. They are purely theoretical objects to study the properties of estimators of the polyserial model when that model is partially misspecified due to data contamination.
While no assumption is made on the specific value of the contamination fraction $\varepsilon$, we impose the restriction $\varepsilon\in[0,0.5)$, which is common in robust statistics (e.g., Hampel et al.,, 1986, p. 67). This restriction ensures proper identification: If half or more of the data points were contaminated, it would no longer be possible to distinguish between contamination and observations generated by the polyserial model, at least not without further assumptions. Under such additional assumptions, also values of $\varepsilon\geq 0.5$ could be considered. We refer to Welz et al., 2025b for a more detailed discussion.

### v2

@@ Abstract
The association between a continuous and an ordinal variable is commonly modeled through the polyserial correlation model. However, this model, which is based on a partially-latent normality assumption, may be misspecified in practice, due to, for example (but not limited to), outliers or careless responses. The typically used maximum likelihood (ML) estimator is highly susceptible to such misspecification: One single observation not generated by partially-latent normality can suffice to produce arbitrarily poor estimates. As a remedy, we propose a novel estimator of the polyserial correlation model designed to be robust against the adverse effects of observations discrepant to that model. The estimator leverages density power divergence estimation to achieve robustness by implicitly downweighting such observations; the ensuing weights constitute a useful tool for pinpointing potential sources of model misspecification. The proposed estimator generalizes ML and is consistent as well as asymptotically Gaussian. As price for robustness, some efficiency must be sacrificed, but substantial robustness can be gained while maintaining more than 98% of ML efficiency. We demonstrate our estimator’s robustness and practical usefulness in simulation experiments and an empirical application in personality psychology where our estimator helps identify outliers. Finally, the proposed methodology is implemented in free open-source software.
@@ 1 Introduction
The polyserial correlation model postulates the existence of a latent continuous variable that underlies and governs the observed ordinal variable through an unobserved discretization process. The correlation between the observed continuous variable and the latent variable is called polyserial correlation, whereas the correlation between the observed continuous and the observed ordinal variable is called point polyserial correlation, where the latter can be computed from the former. For identification, the polyserial correlation model assumes that the observed continuous variable and the latent continuous variable are jointly normally distributed, thereby making it a partially-latent normality model. Estimation is typically conducted by means of maximum likelihood (Cox,, 1974; Olsson et al.,, 1982). However, the validity of the partially-latent normality assumption is often questionable in practice (e.g., Barbiero,, 2025; Demirtas & Hedeker,, 2016; Bedrick,, 1995). Disquietingly, though, central statistical properties of polyserial correlation crucially depend on this very assumption to hold true: Violations of partially-latent normality have been shown to have potentially devastating effects on identification (e.g., Moss & Grønneberg,, 2023) and maximum likelihood estimation (e.g., Bedrick,, 1995) of polyserial correlation. Nonnormality of (partially) latent variables can also introduce serious biases in SEM analyses conducted with software that assumes such normality (e.g., Foldnes & Grønneberg,, 2022, 2020), such as lavaan (Rosseel,, 2012), LISREL (Jöreskog & Sörbom,, 1996), and Mplus (Muthén & Muthén,, 2026).
@@
Motivated by the susceptibility of polyserial correlation to partially-latent nonnormality, the contributions of this paper are twofold. First, we study estimation of the polyserial correlation model when a (possibly empty) subset of the observed data are of low-quality and therefore may have not been generated by a normal distribution, such as (but not limited to) outliers in the continuous variable and/or careless responses in the ordinal variable. Consequently, such observations are uninformative for estimating the polyserial correlation model. This situation is called partial model misspecification because the assumption of partially-normality might be violated for parts of the data, but is satisfied for the remaining data. We demonstrate that already one single uninformative observation can suffice for maximum likelihood (ML) estimation of polyserial correlation to yield arbitrary results and, furthermore, that the accuracy of ML estimation is highly susceptible to even minor misspecification of partially-latent normality. Partial misspecification stems from classic literature on robust statistics (e.g., Huber & Ronchetti,, 2009) where it is known as Huber contamination model, owing to Huber, (1964).
Second, in wake of the non-robustness of ML estimation to uninformative observations generated by partially-latent nonnormality, we propose an alternative estimator that is designed to be robust against such partial misspecification. The proposed methodology applies density power divergence estimation (Basu et al.,, 1998) to the polyserial correlation model and achieves robustness by implicitly downweighting observations that cannot be sufficiently well fitted by that model. The ensuing weights are a useful tool for pinpointing potential sources of (partial) model misspecification. As additional methodological contribution, we devise a simple rescaling of the weights to ensure that they are contained in the unit interval. Overall, to the best of our knowledge, the proposed methodology is the first contamination-robust approach to polyserial correlation.
@@
In line with the partial misspecification framework, the robust estimator allows the model to be misspecified for an unknown fraction of uninformative observations, but makes no assumption on how and where partial misspecification occurs (which may be absent altogether). Consequently, partial misspecification can manifest through an unlimited and unrestricted variety of ways, such as (but not limited to) outliers or careless responses. Conversely, if the polyserial correlation model is correctly specified for all observations in a sample, then the robust estimator is, just like ML, consistent for the true parameter vector, and, therefore, generalizes ML estimation.
Studying their respective behavior under model misspecification, we show that both ML and the robust estimator still converge in probability, but to different limits. Crucially, the robust estimator converges to a parameter vector that is closer to the true parameter vector than ML, thereby gaining its robustness. Moreover, the robust estimator and ML remain asymptotically normally distributed, allowing for statistical inference both under correct and incorrect specification of the polyserial correlation model.
@@
In robust statistics, there is a well-established fundamental tension between robustness and efficiency for estimation procedures for data involving continuous variables (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986). Our estimator is no exception: The price to pay for its enhanced robustness manifests in the form of diminished efficiency compared to ML. However, we show that sacrificing as little as 2% of efficiency suffices to obtain a substantial gain in robustness, so the efficiency loss is only comparatively minor. An additional price of our robust estimator is increased computational intensity. Nevertheless, using our implementation, the robust estimator usually executes in less than two seconds on a regular laptop, so the overall computational burden should remain small in practical applications. This implementation of our proposed methodology is publicly and freely available as part of the R package robcat (for “ROBust CATegorical data analysis”; Welz et al.,, 2025) on CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat.
@@ 2 Literature
Moss & Grønneberg, (2023, Section 2) and Grønneberg et al., (2020, Section 2.5) use partial identification analyses to study distributional misspecification of the polyserial model from a theoretical perspective. Specifically, given known marginal distributions of the observed continuous and latent continuous variables but keeping their joint distribution unspecified, they derive partial identification sets for the partially-latent correlation. They show that the partial identification sets tend to be uninformatively wide. Consequently, polyserial correlation is very sensitive to partially-latent joint nonnormality because different nonnormal distributions can yield widely different correlations. To reduce the width of the partial identification sets, one must make restrictive assumptions on the partially-latent joint distribution, such as Gaussian-like characteristics or equipping it with a parametric structure.
@@
In a broader context, violations of normality assumptions have been studied extensively in the psychometric literature, especially with respect to the distribution of latent variables (e.g., Asparouhov & Muthén,, 2016; Lyhagen & Ornstein,, 2023; Monroe,, 2018; Moss & Grønneberg,, 2023; Roscino & Pollice,, 2006; Yuan et al.,, 2004, and references therein). A particular focus of recent literature has been the polychoric correlation model of Pearson & Pearson, (1922), which models through a latent bivariate normality distribution the association of two latent variables that govern two observed ordinal variables (see Olsson,, 1979, for a modern exposition). Foldnes & Grønneberg, (2020) and Jin & Yang-Wallentin, (2017) show that ML estimation of polychoric correlation is highly susceptible to distributional misspecification, leading to possibly large biases in SEM analyses based on polychoric correlation (Foldnes & Grønneberg,, 2022; Grønneberg & Foldnes,, 2022). [^1] It should be noted that ML estimation of polychoric correlation was originally believed to be fairly robust against distributional misspecification of latent normality (e.g., Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004), an assessment based on simulation experiments using the method of Vale & Maurelli, (1983) for generating nonnormal bivariate data that are then discretized to be ordinal. However, Grønneberg & Foldnes, (2019) show that ordinal data generated in this way are indistinguishable from ordinal data generated by latent normality. Using a method that ensures proper violation of latent normality (the VITA method of Grønneberg & Foldnes,, 2017) reveals strong susceptibility of ML estimation of polychoric correlation to such violations (e.g., Foldnes & Grønneberg,, 2020, 2022). Welz et al., (2026) are concerned with partial misspecification of the polychoric correlation model, where latent normality is violated due to a (possibly empty) subset of data points that were generated by an unspecified and unknown nonnormal process. We use the same (partial) misspecification framework in this paper. Welz et al., (2026) show that having in a sample already about 5% of such observations, commonly referred to as contamination, can suffice for a substantial estimation bias. Contamination might arise due to, for instance but not limited to, careless responding to polytomous items, which has been identified as a major threat to the validity of psychometric analyses (e.g., Alfons & Welz,, 2024; Ward & Meade,, 2023; Arias et al.,, 2020; Bowling et al.,, 2016; Huang et al.,, 2015; Meade & Craig,, 2012, and references therein). As a remedy, Welz et al., (2026) propose a fully efficient contamination-robust estimator of the polychoric correlation model that makes no assumptions on the prevalence and type of contamination (which is possibly absent altogether). Their estimator exploits the theory of $C$-estimation (Welz,, 2024), being a general framework of robust estimation with categorical data.
@@ 3.1 The polyserial correlation model
where we adopt the conventions $\tau_{0}=$-$\infty$ and $\tau_{r}=+\infty$. [^2] Being a random vector comprising a continuous and discrete variable, $(X,Y)$ technically does not have a joint density in a regular sense, i.e., a Randon-Nikodým derivative with respect to the two-dimensional Lebesgue measure. The formally correct way to refer to $p_{XY}$ in (3) would be that it is the Randon-Nikodým derivative of the distribution $P_{XY}$ in (4) with respect to the product of the one-dimensional Lebesgue measure and the counting measure. For the sake of brevity, however, we refer to $p_{XY}$ simply as joint density in this paper, acknowledging an abuse of terminology. The joint distribution function of the observed $(X,Y)$ under the polyserial model can now be expressed as
${P}_{XY}\left(x,y;\bm{\theta}\right)=\mathbb{P}_{\bm{\theta}}\left[X\leq x,Y\leq y\right]=\int_{-\infty}^{x}\sum_{w\leq y}{p}_{XY}\left(u,w;\bm{\theta}\right)\textnormal{d}u=\int_{-\infty}^{x}\int_{-\infty}^{\tau_{y}}{p}_{X\eta}\left(u,v;\bm{\theta}\right)\textnormal{d}v\textnormal{d}u,$ (4)
@@ 3.2 Point polyserial correlation
In addition to the correlation between the observed $X$ and the latent $\eta$, one might also be interested in the correlation between $X$ and the observed ordinal $Y$. The correlation between the observed $X$ and $Y$ is known as point polyserial correlation. In order to identify the desired point polyserial correlation $\mathrm{Cor}\left[X,\ Y\right]$, one needs to assign a numerical interpretation to the $r$ answer categories of $Y$, that is, introduce a scoring system. Given a scoring system, the point polyserial correlation coefficient $\widetilde{\rho}=\mathrm{Cor}\left[X,\ Y\right]$ is identified by the polyserial model and can be estimated by using estimates of the polyserial model parameters $\bm{\theta}$. We provide details in Appendix A.1.
@@ 3.3 Maximum likelihood estimation
While it is in principle possible to robustify the two-step approach approach against contamination by using the same estimation theory as in this paper, doing so would result in substantial theoretical and computational drawbacks. We discuss this in detail in Appendix C.4.
Alternative estimators of polyserial correlation have been proposed by Bedrick & Breslin, (1996), Lord, (1963), and Brogden, (1949), with Bedrick, (1992, 1990), Koopman, (1983) and Kraemer, (1981) studying the theoretical properties of the latter two. We discuss these approaches in more detail in Section 4.2, after conceptualizing misspecification of the polyserial model.
@@ 4.1 Partial misspecification of the polyserial model
The polyserial model is misspecified if at least one observation for the continuous-ordinal variable pair $(X,Y)$ has not been generated by the bivariate partially-latent normality model in (2). Akin to Welz et al., (2026), we consider a partial misspecification framework where only a fraction $(1-\varepsilon)$ of observations in a given sample are generated by an underlying normal distribution $P_{X\eta}$ with true parameter $\bm{\theta}_{*}$, whereas a fixed but unknown fraction $\varepsilon$ of observations are generated from some different but unspecified underlying distribution $H_{X\eta}$. [^3] Being a population quantity and not a sample quantity, it is strictly speaking imprecise to call $\varepsilon$ a fraction. It would be more accurate to call $\varepsilon$ a (contamination) population probability with which one samples from the contamination distribution $H_{X\eta}$ instead of the model distribution $P_{X\eta}$. However, to be consistent with robust statistics literature, we refer to $\varepsilon$ as a fraction throughout this paper. Since $H_{X\eta}$ is unspecified, its correlation structure may differ from $P_{X\eta}$ so that observations $(X,Y)$ generated by the underlying $H_{X\eta}$ (after discretization) may be uninformative for the true polyserial correlation coefficient.
@@
Misspecification models of the type in (8) are standard in the robust statistics literature, where they are known as Huber contamination models, owing to pioneering work of Huber, (1964). Following Welz et al., (2026), we therefore adopt terminology from robust statistics and call $\varepsilon$ the contamination fraction, the uninformative $H_{X\eta}$ the contamination distribution (or simply contamination), and $F_{\varepsilon,X\eta}$ the contaminated distribution. In the case of a zero-valued contamination fraction ($\varepsilon=0$), there is no misspecification so that the polyserial model is correctly specified. Indeed, if $\varepsilon=0$, then $F_{0,X\eta}(\cdot,\cdot)={P}_{X\eta}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$ and $F_{0,XY}(\cdot,\cdot)={P}_{XY}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$, so the partial misspecification framework in (8) nests the correctly specified polyserial model.
Neither the contamination fraction $\varepsilon$ nor the contamination distribution $H_{X\eta}$ in (8) are assumed to be known. Thus, both quantities are left completely unspecified in practice, which “means that we are not making any assumptions on the degree, magnitude, or type of contamination (which is possibly absent altogether)” (Welz et al.,, 2026). Hence, the polyserial model may be misspecified due to an unlimited variety of reasons, for instance but not limited to outliers or nonnormality in the continuous $X$ , and/or careless responding or item misunderstanding in the ordinal $Y$. Because $\varepsilon$ and $H_{X\eta}$ are unspecified in practice, the polyserial model distribution ${P}_{XY}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$ remains the distribution of interest: Our only aim is to estimate the parameter $\bm{\theta}_{*}$ of the polyserial model while reducing the impact of potential contamination in the observed data. Consequently, the contaminated distributions $F_{\varepsilon,XY}$ and $F_{\varepsilon,X\eta}$ are never estimated. They are purely theoretical objects to study the properties of estimators of the polyserial model when that model is partially misspecified due to data contamination.
While no assumption is made on the specific value of the contamination fraction $\varepsilon$, we impose the restriction $\varepsilon\in[0,0.5)$, which is common in robust statistics (e.g., Hampel et al.,, 1986, p. 67). This restriction ensures proper identification: If half or more of the data points were contaminated, it would no longer be possible to distinguish between contamination and observations generated by the polyserial model, at least not without further assumptions. Under such additional assumptions, also values of $\varepsilon\geq 0.5$ could be considered. We refer to Welz et al., (2026) for a more detailed discussion.
@@
We stress that there exist nonnormal joint distributions of $(X,\eta)$ that, after discretizing the latent $\eta$ variable with fixed thresholds, result in the same density for the observed $(X,Y)$ as if the partially-latent $(X,\eta)$ were jointly normal. This phenomenon is known as discretize equivalence of latent variables (Foldnes & Grønneberg,, 2019). It follows that there exist contamination distributions $H_{X\eta}$ and contamination fractions $\varepsilon>0$ in (8) under which the contaminated density $f_{\varepsilon,XY}$ of the observed $(X,Y)$ is exactly equal to the polyserial model density in (3) at the true parameter, that is, $f_{\varepsilon,XY}(x,y)={p}_{XY}\left(x,y;\bm{\theta}_{*}\right)$ for all $(x,y)\in\mathbb{R}\times\mathcal{Y}$. Hence, in this scenario, the polyserial model is misspecified, but the misspecification does not have adverse effects because the true polyserial model density of $(X,Y)$ remains unaffected. To avoid cumbersome notation in our analysis, we follow Welz et al., (2026) and assume consequential misspecification throughout this article, that is, $f_{\varepsilon,XY}(x,y)\neq{p}_{XY}\left(x,y;\bm{\theta}_{*}\right)$ for at least one response $y\in\mathcal{Y}$ whenever $\varepsilon>0$, given $x\in\mathbb{R}$. Nevertheless, “it is silently understood that misspecification need not be consequential” (Welz et al.,, 2026), in which case there is no practical problem and both ML and robust estimator are consistent for the true $\bm{\theta}_{*}$.